Be sure to answer all parts. What is the mass in grams of each quantity of lactic acid (C3H6O3, molar mass 90.08 g/mol), the compound responsible for the aching feeling of tired muscles during vigorous exercise? a. 8.5 × 1024 molecules × 10 g b. 6.93 × 1023 molecules
Answer #1

90.08 gram in 1 mole.

In 1 mole , number of molecules = 6.023 * 10^23

So number of molecules in 90.08 gram = 6.023 * 10^23

Number of molecules in 1 gram = 6.023 * 10^23 / 90

= 0.067 * 10^23 = 6.7* 10^21 molecules

The options are not properly given.